OpenStudy (kiruto222):

What is the perimeter of the pentagon? Regular pentagon with each side labeled twelve feet.

OpenStudy (alexandervonhumboldt2):

perimeter is the sum of all sides of the figure

OpenStudy (alexandervonhumboldt2):

so we have \[\huge\bf P=\color{red}{25+25+25+25+25}=?\]

pooja195 (pooja195):

It says 12 feet so... P=12+12+12+12+12=?